Exercise 2-1 Compute a Predetermined Overhead Rate [LO2-1] Harris Fabrics computes its plantwide predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 20,000 direct labor-hours would be required for the periods estimated level of production. The company also estimated $94,000 of fixed manufacturing overhead cost for the coming period and variable manufacturing overhead of $2.00 per direct labor-hour. Harriss actual manufacturing overhead cost for the year was $123.900 and its actual total direct labor was 21,000 hours. Required Compute the companys plantwide predetermined overhead rate for the year. (Round your answer to 2 decimal places.) Predetermined overhead rate per DLHExercise 2-2 Apply Overhead Cost to Jobs [LO2-2] Luthan Company uses a plantwide predetermined overhead rate of $23.40 per direct labor-hour. This predetermined rate was based on a cost formula that estimated $257,400 of total manufacturing overhead cost for an estimated activity level of 11,000 direct labor- hours. The company incurred actual total manufacturing overhead cost of $249,000 and 10,800 total direct labor-hours during the period. Required: Determine the amount of manufacturing overhead cost that would have been applied to all jobs during the period Manufacturing overheadExercise 2-3 Computing Total Job Costs and Unit Product Costs Using a Plantwide Predetermined Overhead Rate [L02-3 Mickley Companys plantwide predetermined overhead rate is $14.00 per direct labor-hour and its direct labor wage rate is $17.00 per hour. The following information pertains to Job A-500: Direct materials $231 Direct labor $153 Required 1. What is the total manufacturing cost assigned to Job A-500? 2. If Job A-500 consists of 40 units, what is the unit product cost for this job? (Round your answer to 2 decimal places.) Total manufacturing cost 2. Unit product cost per unitExercise 2-12 Computing Predetermined Overhead Rates and Job Costs [LO2-1, LO2-2, LO2-3] Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates: Machine-hours required to support estimated production 100,08e Fixed manufacturing overhead cost Variable manufacturing overhead cost per machine-hour $ 3.ee $65e,000 Required 1. Compute the plantwide predetermined overhead rate. 2. During the year, Job 400 was started and completed. The following information was available with respect to this job: Direct materials Direct labor cost Machine-hours used $450 $210 40 Compute the total manufacturing cost assigned to Job 400 3. If Job 400 includes 52 units, what is the unit product cost for this job? 41f Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would t have established for Job 400?
Answer #1

2-1) Plantwide overhead rate = Estimated overhead/estimated labor hour = (94000/20000+2) = 6.70 per labor hour

2-2) Manufacturing overhead applied = 10800*23.40 = $252720

2-3) Total manufacturing cost

Direct material 231
Direct labor 153
Manufacturing overhead (153/17*14) 126
Total manufacturing cost 510

Unit product cost = 510/40 = 12.75 per unit

2-12) a) Plantwide overhead rate = (650000/100000+3)= 9.50 per machine hour

b) Total manufacturing cost = (450+210+40*9.50) = 1040

c) Unit product cost = 1040/52 = 20

d) Selling price per unit = 20*2.2 = 44